If ln(2pi) 2 (2+sqrt(3)), Then number of roots of the equation 4cos(e x ) = 2 x +2 -x

If ln(2pi)2(2+sqrt(3)), Then number of roots of the equation 4cos(ex) = 2x+2-x

Grade:Upto college level

2 Answers

jagdish singh singh
7 years ago
 Original question …...............................................................................................................................If ln(2pi) 2(2+sqrt(3)), Then number of roots of the equation 4cos(ex) = 2x+2-x
jagdish singh singh
7 years ago
$If $\ln(2\pi)<\log_{2}(2+\sqrt{3})\;,$ Then no. of real roots of $4\cos(e^x) = 2^x+2^{-x}.$